Suppose Duke Energy has a monopoly in the market for electricity in North Carolina. If the fixed cost of Duke Energy is given as $200, 000 and the variable cost of producing Q units of electricity is given as ZcMGI7wkbiwAAAABJRU5ErkJggg==, calculate the Total Costs, Marginal Costs and Average costs of Duke Energy?
Answer #1

Given,

Fixed cost = $200, 000

Variable cost = 6Q2

Total cost = Fixed cost + Variable cost = $200, 000 + 6Q2

Marginal Cost = d/dQ (Total Cost) = d/dQ (200, 000 + 6Q2) = 0 + 6*2Q = 12Q

Average Cost = Total Cost/Output = ($200, 000 + 6Q2) /Q = $200,000/Q + 6Q